# FoMo: A Foundation Model for Mobile Traffic Forecasting with Diffusion Model

### v1

@@ 7. Conclusion
In this paper, we propose FoMo, a foundation model with diffusion models for mobile traffic forecasting. To the best of our knowledge, it is the first foundation model in mobile networks. By capturing the temporal, spatial, human dynamics, and geographical features related to mobile traffic, FoMo exhibits robust multi-task adaptability and zero/few-shot learning capability for diverse tasks across multiple cities, which exhibits good universality. Moreover, we identify the scaling properties of FoMo by examining the model performance with diverse parameter scales and data sizes. We deploy FoMo on the Jiutian platform, where it is used to optimize various aspects of network coverage, throughput, and energy consumption based on its accurate traffic forecasts. FoMo offers valuable insights for integrating multi-dimensional data from users, mobile networking and computing, services, and traffic in mobile networks, further advancing the development of large models in the telecommunications field.

### v3

@@ 7. Conclusion
In this paper, we propose UoMo, a universal model with diffusion models for mobile traffic forecasting. To the best of our knowledge, it is the first universal model in mobile networks that simultaneously supports diverse forecasting tasks including short-term/long-term predictions and generation. By capturing the temporal, spatial, human dynamics, and geographical features related to mobile traffic, UoMo exhibits robust multi-task adaptability and zero/few-shot learning capability for diverse tasks across multiple cities, which exhibits good universality. Moreover, we identify the scaling properties of UoMo by examining the model performance with diverse parameter scales and data sizes. We deploy UoMo on the Jiutian platform, where it is used to optimize various aspects of network coverage, throughput, and energy consumption based on its accurate traffic forecasts. Currently, UoMo has been implemented in Nanning, Guangxi Province, China, for real-time traffic forecasting and data streaming. The model has the potential for large-scale deployment and can effectively assist operators in designing pricing strategies and network expansion, which improves both user experience and economic revenue.
Acknowledgement
This research has been supported in part by the National Key Research and Development Program of China under Grant 2023YFB2904801, the National Natural Science Foundation of China under Grants U23B2030, and the China Postdoctoral Science Foundation under Grant 2023M742010.
Appendix A APPENDIX
A.1. Description of dataset
We collected mobile traffic and user data from 9 different cities, covering over 30,000 base stations, with time granularities ranging from 15 minutes to 1 hour. POI data was crawled from the map service and includes 15 categories, such as lifestyle, entertainment, work, and dining, as shown in Table 7.
Table 7. Description of datasets
Dataset Usage Data Mobile Mobile Time
description traffic users granularity
Beijing 5G data, October, 2021, 4000+ BSs $\checkmark$ $\checkmark$ 1 hour
Shanghai 4G data, August, 2014, 5000+ BSs $\checkmark$ $\checkmark$ 1 hour
Nanjing
Model
training
5G data, February to March, 2021, 6000+ BSs $\checkmark$ $\checkmark$ 15 min
Nanchang 5G data, July, 2023, 5000+ BSs $\checkmark$ $\checkmark$ 30 min
Munich shot tests 4G data, 2022, 2500+ grid-data $\checkmark$ 1 hour
A.2. Description of baselines
Statistical models. Historical moving average method (HA) and ARIMA method that integrate autoregression with average moving. Natural language-based model. Time-LLM describes time series features using natural language and uses these descriptions as prompts into a natural language pre-trained model (LLAMA-7B) for forecasting. Tempo designs temporal prompts with trend and seasonal features for pre-trained models (GPT-2) to predict time series. Spatio-temporal-based models. TimeGPT replaces the Feedforward layer in the transformer with a CNN network and is trained on vast spatio-temporal data. Lagllama uses a set of lag indices to capture different periodic correlations in the time series. CSDI is a conditional diffusion model that uses a masking method for time-series data forecasting and imputation. PatchTST decomposes time series into multiple segments and uses transformers for feature extraction. UniST segments spatio-temporal data and fine-tunes the model using geographical proximity and temporal correlations. Dedicated models for mobile traffic forecasting. SpectraGAN converts mobile traffic generation into an image generation problem and utilizes a CNN-based GAN network for traffic forecasting. KEGAN is a hierarchical GAN that utilizes a self-constructed Urban Knowledge Graph (UKG) to explicitly incorporate urban features during the forecasting process. ADAPTIVE leverages the UKG and a BS aligning scheme to transfer mobile traffic knowledge from one city to another. Open-Diff utilizes open contextual data like satellite images, residential count, and POI distribution to generate mobile traffic data.
A.3. Proof of Lemma 1
The forward chain of DDPM gradually adds Gaussian noise $\epsilon\sim N(0,1)$ to the original data as $q(x_{k}|x_{k-1})=N(\sqrt{1-\beta_{k}}x_{k-1},\beta_{k}\textbf{I})$, $\{\beta_{k}\in(0,1),k\in(1,K)\}$ is a set of scheduled noise weight, and the generated noisy data in step k can be calculated by $x_{k}=\sqrt{\hat{\alpha}_{k}}x_{0}+(1-\hat{\alpha}_{k})\epsilon.$ The reversed chain utilizes a denoising network $p_{\theta}$ to recurrently recover $x_{K}$ to original data $x_{0}$ that yields $p_{\theta}(x_{k-1}|x_{k})=N(\mu_{\theta}(x^{k},k),\sigma_{\theta}(x_{k},k)\textbf{I})$. The objective of the diffusion model is essentially to maximize the log-likelihood function of the denoising network $p_{\theta}$ for the initial data $x_{0}$, i.e.,
(14) $L(\theta)=\mathbb{E}_{x_{0}\sim q(x_{0})}\{-logp_{\theta}(x_{0})\}.$
Subsequently, this function is optimized using the Variational Lower Bound (VLB), which can be expressed as:
Taking the expectation on both sides of the above equation and applying Fubini’s theorem (10.1145/3626235, 62), we can derive:
(16) $L(\theta)=$ $\mathbb{E}_{q(x_{0})}\{-logp_{\theta}(x_{0})\}\leq\mathbb{E}_{q(x_{0})}\bigg{\{}\mathbb{E}_{q_{(}x_{1:T}|x_{0})}\{log\frac{q_{(}x_{1:T}|x_{0})}{p_{\theta}(x_{0:T}|x_{0}))}\}\bigg{\}}$
We can minimize the upper bound of $L(\theta)$ by minimizing $L_{vb}$, thereby maximizing the log-likelihood function of $p_{\theta}$. Ho et al. (10.5555/3495724.3496298, 19) proved that $L_{vb}(\theta)$ can be further parameterized by $\mu_{\theta}(x_{k},k)=\alpha_{k}^{-0.5}[x_{k}-\beta_{k}(1-\hat{\alpha}_{k})^{-0.5}\epsilon_{\theta}(x_{k},k)]$, and $\sigma_{\theta}$ can be parameterized as $\sigma_{\theta}(x_{k},k)=\sqrt{(1-\hat{\alpha}_{k-1})/(1-\hat{\alpha}_{k})\beta_{k}}$. The network $p_{\theta}$ can then be optimized by the following objective:
[…]
The objective in equation (17) is fundamentally equivalent to that of InfoNCE in contrastive learning. We use $p_{\theta}$ to represent the probability in the mutual information as $I(e,y)=p_{\theta}(e_{0:K}|y)/p_{\theta}(e_{0:K})$. In this way, the origin InfoNCE loss can be rewritten as:
where $e^{\prime}$ denotes all negative samples. Referencing the parameterization in (NEURIPS2021_958c5305, 2) where $p_{\theta}(x_{k-1}|x_{k})=\sum_{x_{0}\in q}q(x_{k-1|x_{k},x_{0}})p_{\theta}(x_{0}|x_{k})$, the above loss can be further formulated as:
$\ \ =L^{e}_{vb}-logN\sum\nolimits_{e^{\prime}}L^{e^{\prime}}_{vb}$
where symbol $\doteq$ denotes the loss function we used during the model training process and $\lambda$ is a scaling parameter proportional to $logN$.